OpenStudy (baitcaster13):

A vat of milk has spilled on a tile floor. The milk flow can be expressed with the function r(t) = 4t, where t represents time in minutes and r represents how far the milk is spreading. The spilled milk is creating a circular pattern on the tile. The area of the pattern can be expressed as A(r) = πr2. Part A: Find the area of the circle of spilled milk as a function of time, or A[r(t)]. Show your work. (6 points) Part B: How large is the area of spilled milk after 4 minutes? You may use 3.14 to approximate π in this problem. (4 points)

OpenStudy (baitcaster13):

please someone help im runing out of time my life depends i get this question right

OpenStudy (baitcaster13):

i will become fan for help

OpenStudy (dayakar):

r(t) =4t-------(1) A(r) = pi r^2-------(2) A[r(t)] = A[4t] {from (1)} = \[\pi(4t)^{2}\]

OpenStudy (dayakar):

\[=16\pi*t ^{2}\]

OpenStudy (dayakar):

B) A[r(t)] = 16pi t^2 if t=4mi A[r(4)] = 16*3.14* 4^2 [simplify it u will get the answer]